OpenStudy (anonymous):

For the points (0,0) and (x,-4) find x so that the distance between the points is 5.

OpenStudy (anonymous):

\[\sqrt{(x-0)^2+(-4-0)^2}\]

OpenStudy (anonymous):

=5

OpenStudy (anonymous):

25=x^2+16 25-16=x^2 x^2=9 x=3

OpenStudy (bahrom7893):

y answer ur own question? haha..

OpenStudy (anonymous):

just checking it, i realized when i did it and check back it was correct but i already posted it so i said i will finish it

OpenStudy (sgadi):

Hi pirplec16 .. The work done by you is correct but not complete. The answer is \[\pm 3\]

myininaya (myininaya):

what sgad says is true x=3 or x=-3

OpenStudy (anonymous):

yeah i was wondering if i should put +/- three thanks myininaya!

OpenStudy (anonymous):

\[Let A(x _{1};y _{1}) and B(x_{2};y_{2})\] be 2 points.Then,the distance between them is \[\sqrt{(x_{2}-x_{1})^{2}-(y_{2}-y_{1})^{2}}\]